# Constructing Per-Shot Bitrate Ladders using Visual Information Fidelity

### v1

@@ IV-A VIF Feature Sets
Visual Multimethod Assessment Fusion (VMAF) [25] is a widely used full-reference video quality assessment (VQA) model that has demonstrated excellent correlations against human judgments of video quality. VMAF uses a Support Vector regressor (SVR) model to fuse spatial features from VIF [9], the Detail Loss Metric (DLM) [26], and a simple temporal frame-difference feature computed as the average absolute luminance differences between adjacent frames. VIF [9] is a full-reference image quality assessment (IQA) that predicts the information that could ideally be extracted by visual neurons from the reference image relative to the loss of information from distortion. It uses a Gaussian scale mixture image model expressed in the wavelet domain. VIF features are defined as follows:
@@
where $\mathcal{C}=\{\vec{C}_{i}:i\in I\}$ is a random field (RF) representing a subband of the reference image, $\mathcal{S}=\{\vec{S}_{i}:i\in I\}$ is a RF of positive scalars, and $\mathcal{U}=\{\vec{U}_{i}:i\in I\}$ is a Gaussian RF having mean zero and covariance $\mathbf{C}_{U}$. $\vec{C}_{i}$ and $\vec{U}_{i}$ are M-dimensional vectors and $\vec{U}_{i}$ is independent of $\vec{U}_{j}$ when $i\neq j$. =$\mathcal{E}=\{\vec{E}_{i}:i\in I\}$ is the output of the neural model and $\mathcal{N}=\{\vec{N}_{i}:i\in I\}$ models noise and uncertainty in the wavelet domain as a multivariate Gaussian having mean zero and covariance $\mathbf{C}_{N}=\sigma_{n}^{2}\mathbf{I}$. $I(\vec{C}^{N};\vec{E}^{N}|s^{N})$ represents the information that could ideally be extracted from a particular subband in the image. $\vec{s}^{N}$ is a realization of $S^{N}=(S_{1},S_{2},\dots,S_{N})$ on a particular reference image, which can be thought of as model parameters associated with it. Matrix factorization of (5) yields:
@@
The above features $I_{k}$, $I_{k,b}$, and $I_{k,b}^{j}$ are VIF features at each scale, subband, and along the direction of eigenvector, respectively. Since VIF only captures spatial characteristics, similar to VMAF, we also compute the mean absolute luminance component difference between consecutive frames. As additional temporal features, we apply the same VIF features on differences between the luminance components of consecutive frames, as was done in [27]. Table I shows the nine different feature sets used in our experiments, where $F_{i}$ represents the luminance component of the $i^{\text{th}}$ frame of a source video, and $D_{i}=F_{i}-F_{i-1}$ is the difference between luminance components of adjacent frames. We temporally pool each VIF feature by calculating their means.
@@
TABLE I: List of VIF feature sets.
IV-B Low-Level Features
As discussed in Section II, prior models often employ low-level features like gray-level co-occurrence matrices (GLCM), measures of temporal coherence (TC), normalized cross-correlations (NCC) and other content-sensitive features when modeling RQ-curves and predicting optimal ladder parameters. Traditional spatial information (SI) and temporal information (TI) defined by simple spatial or temporal differencing, respectively, have commonly been used to measure video complexity. However, other authors [28], [29], have shown that chrominance features and DCT-based texture energy features proved more effective when modeling RQ characteristics.
Our study incorporates features derived from an array of prior work on video coding, to develop a comprehensive set of low-level features. We utilize GLCM and TC features, as in [10, 11, 12, 13, 14, 15], and SI, TI, CTI, CF, and CI as employed in [28]. We also compute DCT-based texture energy features as introduced in [16, 29], and extended to chroma components in [17].
The authors of [16], devised a bitrate dependent feature that combines spatial and temporal energy features of luminance: $\log_{2}\left[\sqrt{\frac{h_{Y}}{E_{Y}}}\right]+2\log_{2}(b)$, ($E_{Y}$ is spatial energy, $h_{Y}$ is temporal energy, and $b$ is bitrate) which they used to predict VMAF scores. In addition to the aforementioned feature, we compute features extending computation on the same features expressed on chroma components. Table II shows the list of features we deploy in our work where $GLCM$ is gray-level co-occurrence matrix, $Coherence$ is measure of temporal coherence, $Y$ is luminance, $U$ and $V$ are chroma components, $E_{Y}$, $E_{U}$, $E_{V}$ are spatial energy components, $h_{Y}$, $h_{U}$, $h_{V}$ are temporal energy components, $L_{Y}$, $L_{U}$, $L_{V}$ are luminescence components, $b$ is bitrate, and $q$ is quality. $F_{1}$ operation denotes spatial feature pooling and $F_{2}$ denotes temporal feature pooling.
@@ IV-B Low-Level Features
TABLE II: List of low-level features.
V Experiments and Results
@@
We constructed the experimental dataset by compressing all of the videos using all the encoding settings mentioned in Section III. We divided the dataset into non-overlapping training, validation, and test datasets containing 70, 10, and 20 videos respectively. We ensured that there were no video titles shared between these sets to avoid content learning. The videos in the validation and test datasets were considered together during evaluation, thereby yielding a larger sample size, and these videos were not part of the training data.
In all the experiments, we considered bitrates sampled on a logarithmic scale of base 2, and scaled the VMAF scores by a factor of 100, ensuring they fell within the range of [0,1]. We trained multiple regressors, including Extra-Trees, XG-Boost, and RandomForests, on each task in our experiments. Machine learning models were trained separately on each feature set/approach. We found that the Extra-Trees regressor consistently delivered the best results on all our feature sets on almost every task.
A bitrate ladder contains steps of bitrate and the corresponding optimal resolution for encoding at that bitrate step. When constructing the bitrate ladders, we selected the following bitrates (in kbps): 500, 1000, 2000, 3000, 4000, 5000, 6000, 7000, 8000, 9000, 10500, 12000, 15000. Similarly, a quality ladder contains quality steps based on which full-reference quality model is employed and the corresponding optimal resolution for encoding at that quality step. When constructing quality ladders, we selected the following VMAF score increments: 25, 35, 45, 50, 55, 60, 65, 70, 75, 80, 85, 90, 92.5.
@@
TABLE III: Low-level feature sets.
V-A Predicting Cross-Over Bitrates or VMAF using Low-Level Features
Cross-over points between two resolutions are characterized as intersection points on the RQ curves, marking transitions from a lower to a higher resolution that offers superior quality. This intersection is defined either by a pair of QPs [12, 13], one for each resolution, or by the corresponding bitrate to that intersection [15, 18]. These cross-over points are predicted by extracting features from uncompressed videos, and are then used to construct per-shot bitrate ladders.
Similar to bitrate cross-over points, we designed a method using VMAF cross-over points. The VMAF cross-over points are the intersections of the RQ curves at each resolution where the predicted perceptual qualities are the same. Similar to cross-over bitrates, we predict VMAF cross-over points by extracting features from the source videos, then using them to construct per-shot quality ladders. Fig. 4 shows an example of rate-quality curves and Pareto-Fronts constructed using exhaustive encoding, cross-over bitrates, and VMAF cross-over points on a video sample in the dataset.
@@ V-A Predicting Cross-Over Bitrates or VMAF using Low-Level Features
We compared our proposed methods against a technique that constructs a bitrate ladder using cross-over bitrates. After selecting six different spatial resolutions, we calculated five cross-over bitrates between adjacent resolutions. We used the low-level feature set $\text{LLF}_{1}$ defined in Table III in our experiments. We also used predicted cross-over bitrates between higher resolutions as additional features when predicting cross-over bitrates between lower resolutions. Since our training dataset contains only 70 videos, to avoid the curse of dimensionality, we used recursive feature elimination (RFE) to select nine features that are used to predict each cross-over bitrate. Later, we use the predicted cross-over bitrates to construct per-shot bitrate ladders. As mentioned earlier, we also learned the perceptual counterpart models, viz., that predict VMAF cross-over points using low-level features. Similarly, we applied RFE on the set of low-level features $\text{LLF}_{1}$, to select nine features predictive of VMAF cross-over points, and we later employ these to construct per-shot quality ladders.
@@
V-B Predicting Quality or Bitrate using Low-Level features
We compared our models against [16], where the authors utilized DCT-based texture energy features to predict the quality of compressed videos. Instead of only using DCT texture energy features, we develop a more comprehensive set of features, including metadata like $\log_{2}(b)$, $\frac{w}{3840}$ and $\frac{h}{3840}$, where $b$, $w$ and $h$ are the bitrates, widths, and heights of the compressed videos, respectively. The expanded list of low-level features ($\text{LLF}_{2}$) is shown in Table III. The quality prediction models learned on these features are used to construct per-shot bitrate ladders. To construct per-shot quality ladders, we developed content-dependent bitrate prediction models. These models employ the more comprehensive set of low-level features $\text{LLF}_{3}$ shown in Table III, along with metadata including $\frac{q}{100}$, $\frac{w}{3840}$, and $\frac{h}{3840}$ where $q$ is the VMAF score of each analyzed compressed video, to predict the bitrates of compressed videos. These experiments help us understand the feasibility of predicting the bitrates or qualities of compressed videos using only low-level features extracted from the original source videos.
Table V shows the performances of quality and bitrate prediction models using only low-level video features. It may be observed that the performances of the quality prediction models was slightly higher than that of bitrate prediction models. We found that calculating aggregate PLCC values across all resolutions yields misleading interpretations of model performance at individual resolutions. Hence, we instead report the PLCC values for each resolution.
@@ V-B Predicting Quality or Bitrate using Low-Level features
V-C Predicting Quality or Bitrate using Metadata
Fixed bitrate ladders [3] are generally designed considering a variety of content-independent video characteristics, network conditions, resolutions, and bitrates. We designed a quality prediction model that aims to predict the quality of compressed videos without using any content-dependent features. This quality prediction model is the best statistical fit to the training data based on $\log_{2}(b)$, $\frac{w}{3840}$, and $\frac{h}{3840}$. Although it appears similar to a fixed bitrate ladder, this model is dependent on the characteristics of the videos in the training dataset and the encoding settings used to compress them. We also designed counterpart bitrate prediction models that only train on $\frac{q}{100}$, $\frac{w}{3840}$, and $\frac{h}{3840}$.
@@ V-C Predicting Quality or Bitrate using Metadata
Quality 0.458 0.521 0.552 0.548 0.5 0.421
Bitrate 0.475 0.52 0.535 0.523 0.48 0.408
V-D Predicting Quality or Bitrate using VIF Features
Each VIF feature set contains features extracted from different scales and subbands of uncompressed source videos. Here we examine the performance of these models built using VIF features under different experimental settings and quality constraints. We also designed integrated models trained on VIF feature sets $\text{VIFF}_{i}$ combined with metadata like $\frac{q}{100}$, $\frac{w}{3840}$, and $\frac{h}{3840}$, to learn to predict the bitrates of compressed videos. Comparing the performances of these methods helps us to understand the accuracy, feasibility, and efficacy of constructing bitrate and quality ladders using VIF feature sets.
Fig. 5 shows the performances of bitrate and quality prediction models for each VIF feature set and resolution using color coding. Over the range of implemented CRF values the performances of the two models were similar. Models trained on $\text{VIFF}_{9}$ yielded slightly better correlations against subjectivity when predicting bitrates and qualities across resolutions. The performances of the best performing models was similar to the performances of models trained on low-level features.

### v2

@@ III-B VIF Feature Sets
Visual Multimethod Assessment Fusion (VMAF) [VMAF] is a widely used full-reference video quality assessment (VQA) model that has demonstrated excellent correlations against human judgments of video quality. VMAF uses a Support Vector regressor (SVR) model to fuse spatial features from VIF [VIF], the Detail Loss Metric (DLM) [DLM], and a simple temporal frame-difference feature computed as the average absolute luminance differences between adjacent frames. VIF [VIF] is a full-reference image quality assessment (IQA) that predicts the information that could ideally be extracted by visual neurons from the reference image relative to the loss of information from distortion. It uses a Gaussian scale mixture image model expressed in the wavelet domain.
Fig. 2 shows the differences between the VIF feature extraction in our model, and VIF quality estimation as used in full-reference video quality assessment engines. The block diagram is similar to a figure in [VIF]. The green and red lines in the block diagram represent the paths we considered and excluded, respectively. Unlike during quality assessment, we only compute the information content of the source video as the mutual information between the input and output of the HVS channel. Fig. 3 shows the VIF feature extraction procedure. VIF feature sets consist of features computed at multiple scales and subbands of frames from the reference video, as well as luminance differences between adjacent frames. VIF features are defined as follows:
@@
where $\mathcal{C}=\{\vec{C}_{i}:i\in I\}$ is a random field (RF) representing a subband of the reference image, $\mathcal{S}=\{\vec{S}_{i}:i\in I\}$ is a RF of positive scalars, and $\mathcal{U}=\{\vec{U}_{i}:i\in I\}$ is a Gaussian RF having mean zero and covariance $\mathbf{C}_{U}$. $\vec{C}_{i}$ and $\vec{U}_{i}$ are M-dimensional vectors and $\vec{U}_{i}$ is independent of $\vec{U}_{j}$ when $i\neq j$. $\mathcal{E}=\{\vec{E}_{i}:i\in I\}$ is the output of the neural model and $\mathcal{N}=\{\vec{N}_{i}:i\in I\}$ models noise and uncertainty in the wavelet domain as a multivariate Gaussian having mean zero and covariance $\mathbf{C}_{N}=\sigma_{n}^{2}\mathbf{I}$. $I(\vec{C}^{N};\vec{E}^{N}|s^{N})$ represents the information that could ideally be extracted from a particular subband in the image. $\vec{s}^{N}$ is a realization of $S^{N}=(S_{1},S_{2},\dots,S_{N})$ on a particular reference image, which can be thought of as model parameters associated with it. Matrix factorization of (5) yields:
@@
The above features $I_{k}$, $I_{k,b}$, and $I_{k,b}^{j}$ are VIF features at each scale, subband, and along the direction of eigenvector, respectively. Since VIF only captures spatial characteristics, similar to VMAF, we also compute the mean absolute luminance component difference between consecutive frames. As additional temporal features, we apply the same VIF features on differences between the luminance components of consecutive frames, as was done in [A-visual-information-fidelity-approach-to-video-quality-assessment]. Table I shows the nine different feature sets used in our experiments, where $F_{i}$ represents the luminance component of the $i^{\text{th}}$ frame of a source video, and $D_{i}=F_{i}-F_{i-1}$ is the difference between luminance components of adjacent frames. We temporally pool each VIF feature by calculating their means. Hence, as depicted in Figure 3, the VIF features can be computed using modules already incorporated within the VMAF framework. Furthermore, since the proposed features are exclusively computed on the source video, they can be calculated without employing both compression and quality estimation.
TABLE I: List of VIF feature sets.
@@
III-C Low-Level Features
TABLE II: List of low-level features.
@@ III-C Low-Level Features
As discussed in Section II, prior models often employ low-level features like gray-level co-occurrence matrices (GLCM) [GLCM], measures of temporal coherence (TC) [TC], normalized cross-correlations (NCC) and other content-sensitive features when modeling RQ-curves and predicting optimal ladder parameters. Traditional spatial information (SI) [SI-TI] and temporal information (TI) [SI-TI] defined by simple spatial or temporal differencing, respectively, have commonly been used to measure video complexity. However, other authors [Towards-Perceptually-Optimized-Compression-Of-User-Generated-Content], [VCA], have shown that chrominance features and DCT-based texture energy features proved more effective when modeling RQ characteristics.
Our study incorporates features derived from an array of prior work on video coding, to develop a comprehensive set of low-level features. We utilize GLCM and TC features, as in [Predicting-Video-Rate-Distortion-Curves-using-Textural-Features, Study-of-compression-statistics-and-prediction-of-rate-distortion-curves-for-video-texture, Content-gnostic-Bitrate-Ladder-Prediction-for-Adaptive-Video-Streaming, Efficient-Bitrate-Ladder-Construction-for-Content-Optimized-Adaptive-Video-Streaming, VMAF-based-Bitrate-Ladder-Estimation-for-Adaptive-Streaming, Benchmarking-Learning-based-Bitrate-Ladder-Prediction-Methods-for-Adaptive-Video-Streaming], and SI, TI, contrast information (CTI) [CTI-CI], chrominance information (CI) [CTI-CI], and colorfulness (CF) [CF] as employed in [Towards-Perceptually-Optimized-Compression-Of-User-Generated-Content]. We also compute DCT-based texture energy features as introduced in [Perceptually-Aware-Per-Title-Encoding-for-Adaptive-Video-Streaming, VCA], and extended to chroma components in [Just-Noticeable-Difference-aware-Per-Scene-Bitrate-laddering-for-Adaptive-Video-Streaming].
The authors of [Perceptually-Aware-Per-Title-Encoding-for-Adaptive-Video-Streaming], devised a bitrate dependent feature that combines spatial and temporal energy features of luminance: $\log_{2}\left[\sqrt{\frac{h_{Y}}{E_{Y}}}\right]+2\log_{2}(b)$, ($E_{Y}$ is spatial energy, $h_{Y}$ is temporal energy, and $b$ is bitrate) which they used to predict VMAF scores. In addition to the aforementioned feature, we compute features extending computation on the same features expressed on chroma components. We also devised quality dependent features similar to bitrate dependent features, but using VMAF scores instead of bitrates. Table II shows the list of features we deploy in our work where $GLCM$ is gray-level co-occurrence matrix, $Coherence$ is measure of temporal coherence, $Y$ is luminance, $U$ and $V$ are chroma components, $E_{Y}$, $E_{U}$, $E_{V}$ are spatial energy components, $h_{Y}$, $h_{U}$, $h_{V}$ are temporal energy components, $L_{Y}$, $L_{U}$, $L_{V}$ are luminescence components, $b$ is bitrate, and $q$ is quality. $F_{1}$ operation denotes spatial feature pooling and $F_{2}$ denotes temporal feature pooling.
Similar to VIF features, low-level features are computed only on the source video, thereby eliminating the need for compression and quality estimation. To demonstrate the efficacy of our comprehensive set of low-level features, we plot the correlations among these features. Figure 4(a) presents a $93\times 93$ image matrix representing the absolute PLCC values among all low-level features, excluding bitrate and quality dependent DCT features. Figure 4(b) shows the same PLCC values, but only displaying correlations having absolute values $\geq 0.5$, while zeroing out the rest. It may observed that the low-level features we use show very little correlation with each other, with cross-correlation values between most of the features below $0.5$. This indicates that the considered features are complementary to each other, and can be used to build a robust model. Similarly, Fig. 5(a) and Fig. 5(b) show the correlations among low-level features and VIF features ($\text{VIFF}_{9}$). It may be observed that the considered set of low-level features is little correlated with VIF features (<0.5<0.5). Hence, we may conclude that the considered set of content-aware low-level features and quality-aware VIF features supply only weakly uncorrelated information. The correlation among VIF features ($\text{VIFF}_{9}$) is also evident from Figures 5(a) and 5(b). Although not as effective as the cross-correlation among low-level features, VIF features exhibit less cross-correlation, particularly between spatial and temporal components.
(a) $|\text{PLCC}|$
(b) $|\text{PLCC}|\geq 0.5$
Figure 4: Correlation among low-level features.
(a) $|\text{PLCC}|$
(b) $|\text{PLCC}|\geq 0.5$
Figure 5: Correlation among low-level features and VIF features ($\text{VIFF}_{9}$).
IV Experiments and Results
IV-A Dataset
[…]
Figure 6: Sample frames from each video taken from the BVT-100 4K dataset [Efficient-Bitrate-Ladder-Construction-for-Content-Optimized-Adaptive-Video-Streaming].
IV-B Experiment Settings
(a) Heat-map of RQ points over 8 resolutions and 31 CRFs.
(b) Box-plot of VMAF vs spatial resolution.
(c) Box-plot of VMAF vs CRFs.
[…]
  • Codec, Preset: libx265, medium
[…]
  • CRFs: 16 to 35 (inclusive), 35 to 41 (inclusive) with a skip of 2
  • Constraints: 15 $\leq$ VMAF $\leq$ 95
@@ IV-B Experiment Settings
We constructed the experimental dataset by compressing all of the videos using all the encoding settings mentioned previously. We divided the dataset into non-overlapping training, validation, and test datasets containing 70, 10, and 20 videos respectively. We ensured that there were no video titles shared between these sets to avoid content learning. The videos in the validation and test datasets were considered together during evaluation, thereby yielding a larger sample size, and these videos were not part of the training data.
TABLE III: Low-level feature sets.
@@
IV-C Predicting Cross-Over Bitrates or VMAF using Low-Level Features
Cross-over points between two resolutions are characterized as intersection points on the RQ curves, marking transitions from a lower to a higher resolution that offers superior quality. This intersection is defined either by a pair of QPs [Content-gnostic-Bitrate-Ladder-Prediction-for-Adaptive-Video-Streaming, Efficient-Bitrate-Ladder-Construction-for-Content-Optimized-Adaptive-Video-Streaming], one for each resolution, or by the corresponding bitrate to that intersection [Benchmarking-Learning-based-Bitrate-Ladder-Prediction-Methods-for-Adaptive-Video-Streaming, Ensemble-Learning-for-Efficient-VVC-Bitrate-Ladder-Prediction]. These cross-over points are predicted by extracting features from uncompressed videos, and are then used to construct per-shot bitrate ladders.
Similar to bitrate cross-over points, we designed a method using VMAF cross-over points. The VMAF cross-over points are the intersections of the RQ curves at each resolution where the predicted perceptual qualities are the same. Similar to cross-over bitrates, we predict VMAF cross-over points by extracting features from the source videos, then using them to construct per-shot quality ladders. Fig. 9(d) shows an example of rate-quality curves and Pareto-Fronts constructed using exhaustive encoding, cross-over bitrates, and VMAF cross-over points on a video sample in the dataset.
@@ IV-C Predicting Cross-Over Bitrates or VMAF using Low-Level Features
We compared our proposed methods against a technique that constructs a bitrate ladder using cross-over bitrates. After selecting six different spatial resolutions, we calculated five cross-over bitrates between adjacent resolutions. We used the low-level feature set $\text{LLF}_{1}$ defined in Table III in our experiments. The considered set of features is greater in number than the features employed in [Benchmarking-Learning-based-Bitrate-Ladder-Prediction-Methods-for-Adaptive-Video-Streaming]. We also used predicted cross-over bitrates between higher resolutions as additional features when predicting cross-over bitrates between lower resolutions. Since our training dataset contains only 70 videos, to avoid the curse of dimensionality, we used recursive feature elimination (RFE) to select nine features that are used to predict each cross-over bitrate. Later, we use the predicted cross-over bitrates to construct per-shot bitrate ladders. As mentioned earlier, we also learned the perceptual counterpart models, viz., that predict VMAF cross-over points using low-level features. Similarly, we applied RFE on the set of low-level features $\text{LLF}_{1}$, to select nine features predictive of VMAF cross-over points, and we later employ these to construct per-shot quality ladders.
@@
IV-D Predicting Quality or Bitrate using Low-Level features
We compared our models against [Perceptually-Aware-Per-Title-Encoding-for-Adaptive-Video-Streaming], where the authors utilized DCT-based texture energy features to predict the quality of compressed videos. Instead of only using DCT texture energy features, we develop a more comprehensive set of features, including metadata like $\log_{2}(b)$, $\frac{w}{3840}$ and $\frac{h}{3840}$, where $b$, $w$ and $h$ are the bitrates, widths, and heights of the compressed videos, respectively. The expanded list of low-level features ($\text{LLF}_{2}$) is shown in Table III. The number of features considered is greater than in [Perceptually-Aware-Per-Title-Encoding-for-Adaptive-Video-Streaming]. The quality prediction models learned on these features are used to construct per-shot bitrate ladders. To construct per-shot quality ladders, we developed content-dependent bitrate prediction models. These models employ the more comprehensive set of low-level features $\text{LLF}_{3}$ shown in Table III, along with metadata including $\frac{q}{100}$, $\frac{w}{3840}$, and $\frac{h}{3840}$ where $q$ is the VMAF score of each analyzed compressed video, to predict the bitrates of compressed videos. These experiments help us understand the feasibility of predicting the bitrates or qualities of compressed videos using only low-level features extracted from the original source videos.
Table V shows the performances of quality and bitrate prediction models using only low-level video features. It may be observed that the performances of the quality prediction models was slightly higher than that of bitrate prediction models. We found that calculating aggregate PLCC values across all resolutions yields misleading interpretations of model performance at individual resolutions. Hence, we instead report the PLCC values for each resolution. Tables XII and XIII report the performances of subsets of the low-level features. It may be observed from these results that the comprehensive set of LLF features $\text{LLF}_{2}$ and $\text{LLF}_{3}$ yielded better performances than did individual subsets, with the exception of the DCT texture energy features. The performances of the models trained on DCT texture energy features was slightly better or similar to the performances of models trained on $\text{LLF}_{2}$ or $\text{LLF}_{3}$ across multiple resolutions.
@@ IV-D Predicting Quality or Bitrate using Low-Level features
IV-E Predicting Quality or Bitrate using Metadata
Fixed bitrate ladders [Fixed-Bitrate-Ladder] are generally designed considering a variety of content-independent video characteristics, network conditions, resolutions, and bitrates. We designed a quality prediction model that aims to predict the quality of compressed videos without using any content-dependent features. This quality prediction model is the best statistical fit to the training data based on $\log_{2}(b)$, $\frac{w}{3840}$, and $\frac{h}{3840}$. Although it appears similar to a fixed bitrate ladder, this model is dependent on the characteristics of the videos in the training dataset and the encoding settings used to compress them. We also designed counterpart bitrate prediction models that only train on $\frac{q}{100}$, $\frac{w}{3840}$, and $\frac{h}{3840}$.
@@ IV-E Predicting Quality or Bitrate using Metadata
Quality 0.458 0.521 0.552 0.548 0.500 0.421
Bitrate 0.475 0.520 0.535 0.523 0.480 0.408
IV-F Predicting Quality or Bitrate using VIF Features
Each VIF feature set contains features extracted from different scales and subbands of uncompressed source videos. Here we examine the performance of these models built using VIF features under different experimental settings and quality constraints. We also designed integrated models trained on VIF feature sets $\text{VIFF}_{i}$ combined with metadata like $\frac{q}{100}$, $\frac{w}{3840}$, and $\frac{h}{3840}$, to learn to predict the bitrates of compressed videos. Comparing the performances of these methods helps us to understand the accuracy, feasibility, and efficacy of constructing bitrate and quality ladders using VIF feature sets. Fig. 10 shows the performances of bitrate and quality prediction models for each VIF feature set and resolution using color coding.
The models trained on VIF features extracted along each eigen vector: $\text{VIFF}_{3}$, $\text{VIFF}_{6}$, and $\text{VIFF}_{9}$, yielded better correlations against true and predicted VMAFs/bitrates, as compared to their counterparts trained on VIF features extracted along scales or subbands. The additional mean absolute luminance difference feature computed between consecutive frames improves the performance of the regressors. Similarly, regressors trained on VIF features extracted from frame differences ($\text{VIFF}_{7}$, $\text{VIFF}_{8}$, and $\text{VIFF}_{9}$) delivered superior correlations, as compared to their counterparts trained on VIF features extracted from the original frames and mean absolute luminance component differences. The models trained on $\text{VIFF}_{9}$ yielded better correlations against true and predicted VMAFs/bitrates across resolutions. Interestingly, although VIF features demonstrated higher cross-correlation (Fig. 5), the performances of the best performing models was higher than the performances of models trained on low-level features (Table V, XII and XIII).